OpenStudy (anonymous):

evaluate the expression log[2](√5:(2))

jhonyy9 (jhonyy9):

do you know formula like the first step log(2) (a/b) = log(2) a - log(2) b so and again that log(2) a^x =xlog(2) a hope that these will help you